(x/4) = 2 + (x-3/3)
Answer by edjones(8007)

x/4=2+((x-3)/3) I assume that this is the problem.
3x=24+(4(x-3)) Multiple each side by the LCM (12) to eliminate fractions.
3x=24+4x-12
-x=12
x=-12
.
Ed